# Directed schemes of ideals and cardinal characteristics, I: the meager additive ideal

### v2

@@ Abstract.
The first author was partially supported by the Israel Science Foundation for by grant 2320/23 (2023-2027); The second author was supported by the Grant-in-Aid for Scientific Research (C) 23K03198, Japan Society for the Promotion of Science; and the first and third authors were supported by the grant No. AP0010, Dirección de Tecnología e Investigación and Oficina de Internacionalización, Institución Universitaria Pascual Bravo
@@ 1. Introduction
Yorioka [Yor02] define ideals $\mathcal{I}_{f}$ on $^{\omega}2$, parametrized by $f\in{}^{\uparrow\omega}\omega$, which have Borel bases and characterize $\mathcal{SN}:=\bigcap_{f\in{}^{\uparrow\omega}\omega}\mathcal{I}_{f}$. A further property of the Yorioka ideals is that $\mathcal{I}_{g}\subseteq\mathcal{I}_{f}$ whenever $f\leq^{*}g$.
These ideals led Yorioka to rediscover and significantly improve Seredyński [Ser89] result about the cofinality of $\mathcal{SN}$, and to show that no inequality between $\operatorname{cof}(\mathcal{SN})$ and $\mathfrak{c}$ can be decided in $\mathrm{ZFC}$. More research about the cofinality have been conducted by the first and second authors [Car22, CM23]. One representative result is the following, whose notation is clarified afterward.
@@
Directed schemes of ideals is a crucial notion when it comes to obtaining information on the cardinals of $\mathcal{I}^{*}$ . In fact, the two first authors [Car22, CM23] used $\vec{\mathcal{S}}$ when studying the cofinality of $\mathcal{SN}$. Moreover, in [CM19], they used these ideals to investigate the rest of the cardinals associated with $\mathcal{SN}$. So one of the aims of this paper is to generalize most of the the results established of the cardinals of $\mathcal{SN}$ via Yorioka ideals in the context of the directed scheme of ideals. For example, Theorem 1.2 is consequence of our following main result, which we prove in Section 2.
@@
Let $\vec{\mathcal{I}}$ be a directed scheme of ideals on $S$. Then
@@
    $\operatorname{minadd}(\vec{\mathcal{I}})\leq\operatorname{add}(\mathcal{I}^{*})$.
@@
    $\operatorname{supcov}(\vec{\mathcal{I}})\leq\operatorname{cov}(\mathcal{I}^{*})\leq\operatorname{cov}\left(\prod_{s\in D}\mathrm{Cv}_{\mathcal{I}_{s}}\right)$.
@@
    $\operatorname{non}(\mathcal{I}^{*})=\operatorname{minnon}(\vec{\mathcal{I}})$.
@@
As in the case of the Yorioka ideals, we greatly expect that the directed scheme $\vec{\mathcal{M}}$ is useful to understand the combinatorics of $\mathcal{M}\mathcal{A}$ and its cardinal characteristics. Initially, we intended to approach the following question stated in [CMR25].
@@
Both questions cannot have positive answers simultaneously because $\operatorname{add}(\mathcal{M}\mathcal{A})\leq\mathfrak{b}$ implies $\operatorname{add}(\mathcal{M}\mathcal{A})=\operatorname{add}(\mathcal{M})$ (considering that $\operatorname{add}(\mathcal{M})=\min\{\mathfrak{b},\operatorname{non}(\mathcal{M}\mathcal{A})\}$, due to Pawlikowsky [Paw85]), while $\mathfrak{b}<\operatorname{non}(\mathcal{M}\mathcal{A})$ is consistent with $\mathrm{ZFC}$ ([Paw85, Thm. 2.4]).
@@
To prove this, we refine Shelah’s characterization of $\mathcal{N}\!\mathcal{A}$ from Theorem 1.7 (1) and show a relevant Tukey-connection that gives us a suitable upper bound of $\operatorname{cov}(\mathcal{N}\!\mathcal{A})$.
@@ 2. Directed schemes of ideals
Define the relation $\leq^{*}$ on $^{\omega}\omega$ by $x\leq^{*}y$ means $\forall^{\infty}n<\omega\colon x(n)\leq y(n)$. It is clear that $\langle{}^{\omega}\omega\leq^{*}\rangle$ is a directed preorder and we use $^{\omega}\omega$ to denote its relational system. Then $\mathfrak{b}:=\mathfrak{b}({}^{\omega}\omega)$ and $\mathfrak{d}:=\mathfrak{d}({}^{\omega}\omega)$ are the well known bounding number and dominating number, respectively.
@@
The following result is a useful trick to produce Tukey-connections.
@@
For $i\in\{0,1\}$, let $\mathbf{R}_{i}=\langle X_{i},Y_{i},\sqsubset_{i}\rangle$ be a relational system, an let $\mathbf{R}=\langle S,Y,\sqsubset\rangle$ be a relational system such that $S$ is directed and satisfying that, for $s,s^{\prime}\in S$ and $y\in Y$, if $s^{\prime}\sqsubset y$ and $s\leq_{S}s^{\prime}$ then $s\sqsubset y$.
@@ 1. $(\star)$
As a direct consequence of 3.6, not only we have that any $\mathcal{M}_{I}$ is a $\sigma$-ideal but that $\mathcal{M}_{I}$ is Tukey-equivalent to a clean directed preorder.
@@
Although the answer to the previous question is unknown to us, we have a stronger relation on $\mathbb{I}$ from where $\mathcal{M}_{I^{\prime}}\leq_{\mathrm{T}}\mathcal{M}_{I}$ holds. Namely, for $^{\prime}\in\mathbb{I}, I+II\sqsubseteq^{+}I^{\prime}$ means that, for all but finitely many $n<\omega$, $I^{\prime}_{n}$ is the union of some $I_{k}$’s. Notice that $\langle I^{+},\sqsubseteq^{+}\rangle$ is a preorder but not directed, even more, $\operatorname{cof}(\mathbb{I},\sqsubseteq^{+})=\mathfrak{c}$ (see [CMŠ24, Prop. 3.7]).
@@
We just prove (a) because (b) is consequence of (a), 3.9 and 3.11. For $x\in{}^{\omega}2$, $x\in A_{J,y}^{I^{1}}$ iff $\forall^{\infty}n\,\exists k\in J_{n}\colon x(k)=y(k)$, which is equivent to $\forall^{\infty}n\colon x{\restriction}J_{n}\neq y^{\prime}{\restriction}J_{n}$ where $y^{\prime}:=y+1\mod{2}$. This actually shows $A_{J,y}^{I^{1}}=B_{y^{\prime},J}$ (see 3.1). Therefore, $\mathcal{M}_{I^{1}}=\mathcal{M}$. ∎
@@
[…]
We now look at the cardinal characteristics associated with $\vec{\mathcal{M}}$ and its influence on those with $\mathcal{M}\mathcal{A}$ as well. Since $\vec{\mathcal{M}}$ is a directed squeme of ideals and $\mathcal{M}\mathcal{A}=\vec{\mathcal{M}}^{*}$, by Theorem 2.11 and 2.12 we immediately obtain:
@@
Other relevant result is the following.
@@
In terms of Tukey-connections, $^{\omega}\omega\times\bigvee_{J\in\mathbb{I}}\mathrm{Cv}_{\mathcal{M}_{J}}\leq_{\mathrm{T}}\mathcal{M}_{I}\leq_{\mathrm{T}}\mathcal{M}\leq_{\mathrm{T}}\left(\mathbb{I};\bigvee_{J\in\mathbb{I}}\mathrm{Cv}_{\mathcal{M}_{J}}\right)$.
@@
Thanks to Theorem 2.6, it is enough to prove the Tukey-connections. Since we already have $\mathcal{M}_{I}\leq_{\mathrm{T}}\mathcal{M}$ by 3.12, the following inequalities remain.
@@
For any dominating $D\subseteq\mathbb{I}$, $\mathcal{M}\mathcal{A}\leq_{\mathrm{T}}\prod_{I\in D}\mathcal{M}_{I}\leq_{\mathrm{T}}\mathcal{M}^{D}$. In particular, $\mathcal{M}\mathcal{A}\leq_{\mathrm{T}}\mathcal{M}^{\mathfrak{d}}\leq_{\mathrm{T}}\mathrm{Cv}_{[\operatorname{cof}(\mathcal{M})]^{<\operatorname{add}(\mathcal{M})}}^{\mathfrak{d}}$, $\operatorname{add}(\mathcal{M})\leq\operatorname{add}(\mathcal{M}\mathcal{A})$ and $\operatorname{cof}(\mathcal{M}\mathcal{A})\leq\mathfrak{d}\left(\mathcal{M}^{\mathfrak{d}}\right)\leq\operatorname{cov}\left(([\operatorname{cof}(\mathcal{M})]^{<\operatorname{add}(\mathcal{M})})^{\mathfrak{d}}\right)$.
@@
In [CMR25] we have proved that $\mathbf{R}_{b}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}\mathcal{A}}\leq_{\mathrm{T}}\prod_{b\in D}\mathbf{R}_{b}$ for any $b\in{}^{\omega}\omega$ and any dominating $D\subseteq{}^{\omega}\omega$, and also that $\mathfrak{d}(\mathbf{R}_{b})\leq\operatorname{cof}(\mathcal{M})$. These facts can be deduced from the following results that connects $\mathbf{R}_{b}$ and $\mathrm{Cv}_{\mathcal{M}_{I}}$.
@@
$h\colon\mathbb{I}\times 2^{I}$ $\longmapsto\mathbb{I}_{I}\times{}^{\omega}2$
@@ 4. Consistency results
Thanks to the above, we have the following convenient Tukey-connection.
@@
Here, $\mathfrak{d}_{\theta}$ is the canonical dominating number of $^{\theta}\theta$, which coindices with $\mathfrak{d}(\langle\theta,\theta,\leq\rangle^{\theta})$. It is easy to force $\mathfrak{d}_{\theta}=\nu$, e.g. with $\mathrm{Fn}_{<\theta}(\nu,2)$ (by assuming $\theta^{<\theta}=\theta$ and $\nu^{\theta}=\nu$ in the ground model).

### v3

@@ Abstract.
The first author was partially supported by the Israel Science Foundation by grant 2320/23 (2023-2027); The second author was supported by the Grant-in-Aid for Scientific Research (C) 23K03198, Japan Society for the Promotion of Science; and the first and third authors were supported by the grant No. AP0010, Dirección de Tecnología e Investigación and Oficina de Internacionalización, Institución Universitaria Pascual Bravo
@@ 1. Introduction
Yorioka [Yor02] defined ideals $\mathcal{I}_{f}$ on $^{\omega}2$, parametrized by $f\in{}^{\uparrow\omega}\omega$, which have Borel bases and characterize $\mathcal{SN}:=\bigcap_{f\in{}^{\uparrow\omega}\omega}\mathcal{I}_{f}$. A further property of the Yorioka ideals is that $\mathcal{I}_{g}\subseteq\mathcal{I}_{f}$ whenever $f\leq^{*}g$.
These ideals led Yorioka to rediscover and significantly improve Seredyński [Ser89] result about the cofinality of $\mathcal{SN}$, and to show that no inequality between $\operatorname{cof}(\mathcal{SN})$ and $\mathfrak{c}$ can be decided in $\mathrm{ZFC}$. More research about the cofinality has been conducted by the first and second authors [Car22, CM23]. One representative result is the following, whose notation is clarified afterward.
@@
Directed schemes of ideals is a crucial notion when it comes to obtaining information on the cardinals of $\mathcal{I}^{*}$ . In fact, the two first authors [Car22, CM23] used $\vec{\mathcal{S}}$ when studying the cofinality of $\mathcal{SN}$. Moreover, in [CM19], they used these ideals to investigate the rest of the cardinals associated with $\mathcal{SN}$. So one of the aims of this paper is to generalize most of the results established of the cardinals of $\mathcal{SN}$ via Yorioka ideals in the context of the directed scheme of ideals. For example, Theorem 1.2 is a consequence of our following main result, which we prove in Section 2.
@@
Let $\vec{\mathcal{I}}$ be a directed scheme of ideals on $S$. Then:
@@
    $\operatorname{minadd}(\vec{\mathcal{I}})\leq\operatorname{add}(\mathcal{I}^{*})$;
@@
    $\operatorname{supcov}(\vec{\mathcal{I}})\leq\operatorname{cov}(\mathcal{I}^{*})\leq\operatorname{cov}\left(\prod_{s\in D}\mathrm{Cv}_{\mathcal{I}_{s}}\right)$;
@@
    $\operatorname{non}(\mathcal{I}^{*})=\operatorname{minnon}(\vec{\mathcal{I}})$;
@@
As in the case of the Yorioka ideals, we greatly expect that the directed scheme $\vec{\mathcal{M}}$ is useful to understand the combinatorics of $\mathcal{M}\mathcal{A}$ and its cardinal characteristics. Initially, we intended to approach the following questions stated in [CMR25].
@@
Both questions cannot have positive answers simultaneously because $\operatorname{add}(\mathcal{M}\mathcal{A})\leq\mathfrak{b}$ implies $\operatorname{add}(\mathcal{M}\mathcal{A})=\operatorname{add}(\mathcal{M})$ (considering that $\operatorname{add}(\mathcal{M})=\min\{\mathfrak{b},\operatorname{non}(\mathcal{M}\mathcal{A})\}$, due to Pawlikowski [Paw85]), while $\mathfrak{b}<\operatorname{non}(\mathcal{M}\mathcal{A})$ is consistent with $\mathrm{ZFC}$ ([Paw85, Thm. 2.4]).
@@
To prove this, we refine Shelah’s characterization of $\mathcal{N}\!\mathcal{A}$ from Theorem 1.7 (1) and show a relevant Tukey connection that gives us a suitable upper bound of $\operatorname{cov}(\mathcal{N}\!\mathcal{A})$.
@@ 2. Directed schemes of ideals
Define the relation $\leq^{*}$ on $^{\omega}\omega$ by $x\leq^{*}y$ means $\forall^{\infty}n<\omega\colon x(n)\leq y(n)$. It is clear that $\langle{}^{\omega}\omega\leq^{*}\rangle$ is a directed preorder and we use $^{\omega}\omega$ to denote its relational system. Then $\mathfrak{b}:=\mathfrak{b}({}^{\omega}\omega)$ and $\mathfrak{d}:=\mathfrak{d}({}^{\omega}\omega)$ are the well-known bounding number and dominating number, respectively.
@@
The following result is a useful trick to produce Tukey connections.
@@
For $i\in\{0,1\}$, let $\mathbf{R}_{i}=\langle X_{i},Y_{i},\sqsubset_{i}\rangle$ be a relational system, and let $\mathbf{R}=\langle S,Y,\sqsubset\rangle$ be a relational system such that $S$ is directed and satisfying that, for $s,s^{\prime}\in S$ and $y\in Y$, if $s^{\prime}\sqsubset y$ and $s\leq_{S}s^{\prime}$ then $s\sqsubset y$.
@@ 1. $(\star)$
As a direct consequence of 3.6, not only do we have that any $\mathcal{M}_{I}$ is a $\sigma$-ideal but also that $\mathcal{M}_{I}$ is Tukey-equivalent to a clean directed preorder.
@@
Although the answer to the previous question is unknown to us, we have a stronger relation on $\mathbb{I}$ from where $\mathcal{M}_{I^{\prime}}\leq_{\mathrm{T}}\mathcal{M}_{I}$ holds. Namely, for $^{\prime}\in\mathbb{I}, I+II\sqsubseteq^{+}I^{\prime}$ means that, for all but finitely many $n<\omega$, $I^{\prime}_{n}$ is the union of some $I_{k}$’s. Notice that $\langle I^{+},\sqsubseteq^{+}\rangle$ is a preorder but not directed, even more, $\operatorname{cof}(\mathbb{I},\sqsubseteq^{+})=\mathfrak{c}$ (see [CMŠ26, Prop. 3.7]).
@@
We just prove (a) because (b) is consequence of (a), 3.9 and 3.11. For $x\in{}^{\omega}2$, $x\in A_{J,y}^{I^{1}}$ iff $\forall^{\infty}n\,\exists k\in J_{n}\colon x(k)=y(k)$, which is equivalent to $\forall^{\infty}n\colon x{\restriction}J_{n}\neq y^{\prime}{\restriction}J_{n}$ where $y^{\prime}:=y+1\mod{2}$. This actually shows $A_{J,y}^{I^{1}}=B_{y^{\prime},J}$ (see 3.1). Therefore, $\mathcal{M}_{I^{1}}=\mathcal{M}$. ∎
@@
$X\in\mathcal{M}\mathcal{A}$ $\text{\ iff\ }\forall A\in\mathcal{M}\colon X+A\in\mathcal{M}$
$\text{\ iff\ }\forall(I,y)\in\mathbb{I}\times{}^{\omega}2\,\exists(J,y^{\prime})\in\mathbb{I}\times{}^{\omega}2\colon X+A_{I,y}^{I^{1}}=\underbrace{\bigcup_{x\in X}x+A_{I,y}^{I^{1}}}_{\bigcup_{x\in X}A_{I,x+y}^{I^{1}}}\subseteq A_{J,y^{\prime}}^{I^{1}}$
$\text{\ iff\ }\forall(I,y)\in\mathbb{I}\times{}^{\omega}2\,\exists(J,y^{\prime})\in\mathbb{I}\times{}^{\omega}2\,\forall x\in X\colon A_{I,x+y}^{I^{1}}\subseteq A_{J,y^{\prime}}^{I^{1}}$
$\text{\ iff\ }\forall I\in\mathbb{I}\,\exists(J,y)\in\mathbb{I}\times{}^{\omega}2\,\forall x\in X\colon A_{I,x}^{I^{1}}\subseteq A_{J,y}^{I^{1}}\text{ (using a translation)}$
$\text{\ iff\ }\forall I\in\mathbb{I}\,\exists(J,y)\in\mathbb{I}\times{}^{\omega}2\,\forall x\in X\,\forall^{\infty}n\in\omega\,\exists\ell<\omega\colon I_{\ell}\subseteq J_{n}\,\wedge\,x{\restriction}I_{\ell}=y{\restriction}I_{\ell}$
$\text{\ iff\ }\forall I\in\mathbb{I}\,\exists(J,y)\in\mathbb{I}\times{}^{\omega}2\colon X\subseteq A_{J,y}^{I}$
$\text{\ iff\ }\forall I\in\mathbb{I}\colon X\in\mathcal{M}_{I}$
$\text{\ iff\ }X\in\bigcap_{I\in\mathbb{I}}\mathcal{M}_{I}.\qed$
We now look at the cardinal characteristics associated with $\vec{\mathcal{M}}$ and its influence on those with $\mathcal{M}\mathcal{A}$ as well. Since $\vec{\mathcal{M}}$ is a directed scheme of ideals and $\mathcal{M}\mathcal{A}=\vec{\mathcal{M}}^{*}$, by Theorem 2.11 and 2.12 we immediately obtain:
@@
Another relevant result is the following.
@@
In terms of Tukey connections, $^{\omega}\omega\times\bigvee_{J\in\mathbb{I}}\mathrm{Cv}_{\mathcal{M}_{J}}\leq_{\mathrm{T}}\mathcal{M}_{I}\leq_{\mathrm{T}}\mathcal{M}\leq_{\mathrm{T}}\left(\mathbb{I};\bigvee_{J\in\mathbb{I}}\mathrm{Cv}_{\mathcal{M}_{J}}\right)$.
@@
Thanks to Theorem 2.6, it is enough to prove the Tukey connections. Since we already have $\mathcal{M}_{I}\leq_{\mathrm{T}}\mathcal{M}$ by 3.12, the following inequalities remain.
@@
For any dominating $D\subseteq\mathbb{I}$,
$\mathcal{M}\mathcal{A}\leq_{\mathrm{T}}\prod_{I\in D}\mathcal{M}_{I}\leq_{\mathrm{T}}\mathcal{M}^{D}.$
In particular, $\mathcal{M}\mathcal{A}\leq_{\mathrm{T}}\mathcal{M}^{\mathfrak{d}}\leq_{\mathrm{T}}\mathrm{Cv}_{[\operatorname{cof}(\mathcal{M})]^{<\operatorname{add}(\mathcal{M})}}^{\mathfrak{d}}$, $\operatorname{add}(\mathcal{M})\leq\operatorname{add}(\mathcal{M}\mathcal{A})$ and $\operatorname{cof}(\mathcal{M}\mathcal{A})\leq\mathfrak{d}\left(\mathcal{M}^{\mathfrak{d}}\right)\leq\operatorname{cov}\left(([\operatorname{cof}(\mathcal{M})]^{<\operatorname{add}(\mathcal{M})})^{\mathfrak{d}}\right)$.
@@
In [CMR25] we have proved that $\mathbf{R}_{b}\leq_{\mathrm{T}}\mathrm{Cv}_{\mathcal{M}\mathcal{A}}\leq_{\mathrm{T}}\prod_{b\in D}\mathbf{R}_{b}$ for any $b\in{}^{\omega}\omega$ and any dominating $D\subseteq{}^{\omega}\omega$, and also that $\mathfrak{d}(\mathbf{R}_{b})\leq\operatorname{cof}(\mathcal{M})$. These facts can be deduced from the following result that connects $\mathbf{R}_{b}$ and $\mathrm{Cv}_{\mathcal{M}_{I}}$.
@@
$h\colon\mathbb{I}\times 2^{I}$ $\longrightarrow\mathbb{I}_{I}\times{}^{\omega}2$
@@ 4. Consistency results
Thanks to the above, we have the following convenient Tukey connection.
@@
Here, $\mathfrak{d}_{\theta}$ is the canonical dominating number of $^{\theta}\theta$, which coincides with $\mathfrak{d}(\langle\theta,\theta,\leq\rangle^{\theta})$. It is easy to force $\mathfrak{d}_{\theta}=\nu$, e.g. with $\mathrm{Fn}_{<\theta}(\nu,2)$ (by assuming $\theta^{<\theta}=\theta$ and $\nu^{\theta}=\nu$ in the ground model).